OpenStudy (anonymous):

write the point slope form of the equation of the line through the given point with the given slope

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when given m = slope and point (x1, y1)

jigglypuff314 (jigglypuff314):

"write the point slope form of the equation of the line through the given point with the given slope through 5,0 slope = -2" so plug in m = -2 x1=5 and y1=0 into y - (y1) = (m) (x - (x1) )

OpenStudy (anonymous):

thanks

jigglypuff314 (jigglypuff314):

you're welcome :)